### v3

# Mantis: A Simulation-Grounded Foundation Model for Disease Forecasting

Abstract
Infectious disease forecasting in novel outbreaks or low-resource settings is hampered by the need for disease-specific data, bespoke training, and expert tuning. We introduce Mantis, a foundation model trained entirely on mechanistic simulations, which enables out-of-the-box forecasting across diseases, regions, and outcomes, even in settings with limited historical data. We evaluated Mantis against 48 forecasting models across six diseases with diverse transmission modes, assessing both point forecast accuracy (mean absolute error) and probabilistic performance (weighted interval score and coverage). Despite using no real-world data during training, Mantis achieved lower mean absolute error than all models in the CDC’s COVID-19 Forecast Hub when backtested on early pandemic forecasts. Across all other diseases tested, including respiratory, vector-borne, and waterborne pathogens, Mantis consistently ranked in the top two models across all evaluation metrics. Notably, Mantis generalized to diseases with transmission mechanisms not represented in its training data, demonstrating that it captures fundamental contagion dynamics rather than memorizing disease-specific patterns. These capabilities position Mantis as a practical foundation for disease forecasting: general-purpose, accurate, and deployable where traditional models fail.
Introduction
Accurate forecasting of infectious disease dynamics is fundamental to effective public health response. During the COVID-19 pandemic, forecasts touched nearly every aspect of daily life—from how scarce hospital beds were allocated to if restaurants could open and whether families could send their children to school [1]. Yet despite their importance, disease forecasts often fail in practice, particularly when they are most needed. During the first two years of the COVID-19 pandemic, the performance of many models in the CDC Forecast Hub was worse than a naïve baseline that simply projected the most recent value forward [2]. This highlights a key limitation of many operational systems: they struggle to provide accurate, actionable forecasts precisely when uncertainty is highest and decisions are most urgent.
Several structural limitations contribute to this breakdown. First, most models require large volumes of disease-specific historical data, which are rarely available early in outbreaks. Second, they rely on extensive manual tuning by expert teams, limiting scalability and access in under-resourced health departments, such as those in rural counties or low- and middle-income countries. Third, they generalize poorly across diseases, locations, or targets, necessitating new models for every context. In addition, their forecasting horizons are typically short: most models only predict 1–4 weeks ahead [2, 3, 4], far too short-term for real-world public health planning. Key decisions, such as ramping up hospital capacity, acquiring ventilators, or coordinating vaccination campaigns, often require 6–8 weeks of lead time. Without reliable long-range forecasts, these decisions are either delayed or made reactively, often at considerable cost [5].
Typically, projections beyond 4 weeks are restricted to scenario analysis, which explores hypothetical outcomes conditional on specific policy interventions or behavioral changes [6, 7]. However, unconditional forecasts at longer horizons remain valuable for public health planning. We define forecasts as unconditional projections of expected outcomes, versus scenario analyses that specify hypothetical interventions. By this definition, temporal horizon does not determine whether a projection is a forecast or a scenario—an 8-week unconditional projection is a forecast, while a 2-week conditional projection is still scenario analysis. Public health decisions such as hospital capacity expansion and vaccination campaigns often require 6-8 weeks of lead time, motivating the development of models capable of reliable longer-range forecasting without any policy assumptions.
Addressing these challenges (limited historical data, poor cross-disease generalization, restricted forecast horizons, and the need for extensive expert tuning) requires a different approach to model training. Simulation-Grounded Neural Networks (SGNNs) provide such a framework [8]. SGNNs learn entirely from large-scale, highly realistic mechanistic simulations, using data generated by domain-specific simulators to train flexible neural networks. In epidemiology, this means generating synthetic outbreaks from mechanistic disease models that capture processes such as transmission, progression, immunity, seasonality, and intervention effects, then training models on these simulated trajectories so they can forecast real-world outbreaks they have never encountered before. By grounding learning in simulated systems that encode known mechanisms—whether physical, biological, or social—SGNNs enable models to generalize to real-world settings where data are limited, noisy, or missing.
We present Mantis, the first large-scale foundation model built on the SGNN paradigm. Building on early demonstrations of simulation-grounded forecasting [9], Mantis is trained entirely on mechanistic simulations without access to any real-world training data. It learns from over 400 million simulated days of outbreak dynamics spanning diverse pathogens, transmission routes (human-to-human, vector-borne, environmental), intervention strategies, immunity profiles, and surveillance artifacts such as underreporting and delays. These simulations encode core epidemiological mechanisms, allowing Mantis to internalize generalizable contagion dynamics rather than extrapolating from historical trends. Unlike traditional models, Mantis requires no disease-specific tuning or retraining, yet it produces accurate probabilistic forecasts out-of-the-box across diseases, geographies, and targets (Figure 1).
Figure 1: Conceptual overview of Mantis. Mantis is a simulation-grounded foundation model trained entirely on synthetic outbreaks generated by mechanistic epidemiological models. The training pipeline begins with a modular simulator that encodes diverse outbreak mechanisms, including multiple transmission modes (human-to-human, vectorborne, environmental), progression dynamics, intervention strategies, and population structures. These simulations are then passed through an observation model that incorporates real-world surveillance effects such as underreporting, stochasticity, and reporting delays. Trained on over 400 million simulated days, Mantis forecasts directly from real-world time series at inference time. Key advantages include accurate out-of-the-box performance, long-range forecasting capability, and mechanistic interpretability via back-to-simulation attribution.
We evaluate Mantis against 48 models across forecasting tasks spanning six real-world diseases: seasonal influenza, early COVID-19, dengue, hepatitis B, smallpox, and scarlet fever. These evaluations cover multiple transmission modes, outbreak phases, surveillance systems, and historical eras. We assessed both point forecast accuracy and probabilistic performance. When benchmarked against CDC Forecast Hub models on early pandemic COVID-19 forecasts, Mantis achieved the lowest mean absolute error despite using no COVID-specific data or covariates in training and was tied for the second lowest weighted interval score. Across all other diseases—including those with transmission mechanisms absent from its training data—Mantis consistently ranked among the top two models on all evaluation metrics, demonstrating robust generalization to diverse epidemiological settings.
These capabilities establish Mantis as a general-purpose forecasting system offering a scalable solution for modeling in data-scarce or rapidly evolving settings. To support deployment, we release Mantis as an open-source repository. Users can generate forecasts from a single historical time series without any machine learning expertise. 222github.com/carsondudley1/Mantis: Pretrained models, example data, and an interactive forecasting demo are available. Mantis is currently contributing forecasts to a new local-scale forecasting hub within Michigan for the 2025-2026 respiratory virus season and will be submitted to the CDC Forecast Hub to support real-time public health decision-making.
Results
We assessed Mantis’s forecasting accuracy across a set of real-world forecasting tasks spanning six diseases: seasonal influenza [10, 11], early COVID-19 [12], dengue [13], hepatitis B [14], smallpox [15], and scarlet fever [16]. For each task, we benchmarked Mantis against expert-tuned statistical and machine learning models, including published baselines and submissions to the CDC COVID-19 Forecast Hub [2] (see Methods). We evaluated both point forecast accuracy using mean absolute error (MAE) and probabilistic forecasting performance using weighted interval score (WIS) and coverage metrics (e.g., how often did the 95% CI actually contain the true value).
Mantis Achieves Competitive Performance Across Diseases and Tasks
For each disease, we evaluated performance across multiple spatial and temporal resolutions, including U.S. states and smaller regions within states. We evaluated both point forecast performance (mean absolute error, MAE) and probabilistic metrics (weighted interval score, WIS; 50% and 95% coverage). WIS is the standard probabilistic evaluation metric for disease forecasting [17], providing a comprehensive assessment that accounts for both forecast accuracy and uncertainty calibration across predicted quantiles. Following the evaluation framework established in [2], we report relative MAE and relative WIS, where metrics are normalized by the performance of a naïve baseline that forecasts the current value forward with uncertainty intervals derived from historical variance (see Methods). Relative metrics offer two key advantages over absolute error: they provide an interpretable measure of the information contributed by a model beyond simple persistence, and they enable meaningful comparisons across different forecast targets and geographic scales, where absolute case or death counts may vary by orders of magnitude.
We compare Mantis against a range of published and re-implemented models, including top submissions to the CDC COVID-19 Forecast Hub, statistical baselines (exponential smoothing and SARIMA), LSTM neural networks retrained at each forecast point, and Chronos, a state-of-the-art time-series foundation model [18].
Mantis achieved competitive performance across diseases and metrics (Table 1). For early pandemic COVID-19 forecasts at the state level, Mantis achieved the lowest MAE among evaluated models, including the CDC Forecast Hub ensemble that combines dozens of expert-tuned submissions. The CDC ensemble achieved slightly better WIS (0.61 vs 0.63), while UMass-MechBayes achieved marginally better 95% coverage (0.93 vs 0.91), demonstrating that multiple approaches can excel on different aspects of probabilistic forecasting. Although results have shown that some general time series forecasting foundation models can confer reasonable results for infectious disease forecasting [19], we did not find that any we tested gave competitive performance with Mantis or our other baseline models, so we did not include them in our analysis.
Across the five additional diseases tested—influenza hospitalizations, dengue, smallpox, scarlet fever, and hepatitis B—Mantis consistently ranked first or second on both point forecast accuracy (MAE) and probabilistic performance (WIS).
Example forecast trajectories illustrate Mantis’s ability to match the timing and magnitude of real-world outbreaks across diverse epidemiological settings (Figure 3). Forecasts include uncertainty intervals derived from the model’s predicted quantiles, enabling direct use in probabilistic planning. Because Mantis is trained entirely on broad mechanistic simulations rather than historical observations, it requires no model fitting or fine-tuning for new targets.
Table 1: Forecasting results across diseases and models. For each metric, the best model is highlighted in green and bold, and the second-best is blue and underlined.
Disease Model
Relative MAE
Relative WIS
50% coverage
95% coverage
State-level COVID Mantis 0.65 0.63 0.51 0.91
CDC ensemble 0.66 0.61 0.53 0.90
UMass-MechBayes 0.67 0.63 0.56 0.93
GT-DeepCOVID 0.82 0.75 0.41 0.84
UCSD_NEU-DeepGLEAM 0.78 0.83 0.70 0.91
SARIMA 1.08 1.10 0.44 0.81
LSTM 1.37 1.54 0.33 0.70
Local-scale flu hosp. Mantis 0.85 0.86 0.51 0.92
ETS 1.15 1.32 0.39 0.78
LSTM 1.21 1.33 0.40 0.79
SARIMA 1.10 1.00 0.47 0.78
Dengue cases Mantis 0.75 0.80 0.47 0.90
ETS 0.99 1.04 0.45 0.90
LSTM 1.34 1.62 0.37 0.73
SARIMA 0.94 1.58 0.73 0.92
Smallpox cases Mantis 0.83 0.84 0.45 0.87
ETS 1.00 1.06 0.44 0.86
LSTM 1.31 1.52 0.33 0.71
SARIMA 0.95 1.04 0.53 0.88
ILI syndromic Mantis 0.83 0.88 0.46 0.89
ETS 1.00 1.05 0.43 0.83
LSTM 1.19 1.30 0.41 0.80
SARIMA 1.03 0.97 0.45 0.88
Scarlet Fever cases Mantis 0.80 0.81 0.47 0.91
ETS 0.89 0.95 0.46 0.89
LSTM 0.87 1.04 0.36 0.80
SARIMA 1.05 1.08 0.56 0.90
Hepatitis B cases Mantis 0.85 0.84 0.47 0.90
ETS 0.86 0.89 0.46 0.88
LSTM 0.82 1.07 0.37 0.78
SARIMA 0.86 0.84 0.58 0.91
Covariates Improve Forecast Accuracy Across Horizons
We evaluated two variants of Mantis to assess the impact of additional information on forecasting performance: a covariate-free version, which receives only the time series to be forecast (e.g., deaths), and a covariate-aware version, which also receives related signals (e.g., cases or hospitalizations). This design mirrors real-world forecasting settings, where models may have access to upstream indicators with predictive value. (See Methods for training details for each model.)
For example, when forecasting COVID-19 deaths, the covariate-aware Mantis version receives both historical deaths and upstream case data, while the covariate-free version sees only the death time series. Both models use identical architecture and training data; the only difference is the presence or absence of additional input channels.
Across forecast horizons (2, 4, 6, and 8 weeks), the covariate-aware version consistently outperformed its covariate-free counterpart (Figure 2a), achieving lower mean absolute error across diseases and targets. These improvements were particularly pronounced at longer horizons, where upstream indicators (e.g., cases leading to hospitalizations, then deaths) provide early signals that improve downstream predictions.
These results highlight the benefit of incorporating complementary epidemiological signals when available and demonstrate that Mantis can flexibly leverage additional information without retraining. In data-rich settings, covariate-aware forecasts yield improved performance; in low-data contexts, the covariate-free version retains strong out-of-the-box accuracy, making both variants useful in practice depending on context and data availability.
Figure 2: Covariate integration improves accuracy. Mantis maintains calibrated uncertainty across forecast horizons. (a) Including covariates (e.g., using cases to predict hospitalizations) consistently improves Mantis’s accuracy across all forecast horizons. Relative MAE shown for COVID-19 mortality forecasts with (blue) and without (orange) hospitalization covariates across 2, 4, 6, and 8-week horizons. (b) Mantis’s prediction interval widths increase appropriately with forecast horizon, reflecting growing uncertainty in longer-range predictions. The 50% confidence interval width (blue) grows from approximately 50 to 80 deaths, while the 95% interval width (orange) expands from approximately 110 to 180 deaths over 1 to 4 weeks ahead when forecasting COVID-19 mortality. This systematic widening demonstrates the model becoming appropriately less confident as prediction horizon increases.
Mantis Enables Accurate Forecasts at Longer Horizons
We evaluated Mantis’s long-range forecasting performance at an 8-week horizon across the same benchmark tasks and baselines described above. Mantis consistently achieved the lowest relative MAE across all diseases and targets tested, demonstrating its ability to extrapolate epidemiological dynamics further into the future than prior models. At the 8-week horizon, Mantis achieved lower relative error than all models in the CDC Forecast Hub at 4 weeks, and lower raw error than the median 4-week error of Forecast Hub models, demonstrating long-range accuracy that exceeds the short-term limits of existing expert-tuned systems.
Visual examples of 8-week forecasts are shown in figure 3b. Prediction intervals widen with time (figure 2b), maintaining calibrated uncertainty that reflects forecast difficulty (calibration results in table 1).
This capability is not enabled by manual tuning or disease-specific details, but by scale: Mantis learns from millions of simulated outbreak trajectories, encompassing diverse transmission patterns and intervention regimes. Unlike historically trained models, which rely on narrow slices of observed data, simulation-trained models like Mantis can internalize general patterns that support long-range prediction.
Figure 3: Mantis Produces Accurate and Generalizable Forecasts Across Diseases and Geographies. (a) Four-week-ahead forecasts (blue dashed line and shaded 90% CI) compared to observed outcomes (black) for COVID-19 mortality in Minnesota and influenza-like illness (ILI) in Michigan. In the latter, Mantis demonstrates its foundation model capacity by accurately forecasting syndromic inputs despite never being trained with syndromic data. (b) Eight-week-ahead forecasts for four historical outbreaks, highlighting Mantis’s ability to generalize across time, space, and transmission mode. In Alagoas, Brazil, Mantis forecasts the 2004 dengue surge prematurely, highlighting a failure in temporal calibration due to the absence of covariates, but still estimates the peak height with remarkable accuracy despite being 18 weeks early. Forecasts for hepatitis B in Florida—a chronic, bloodborne disease type absent from Mantis’s training set—further demonstrate its capacity to generalize to out-of-distribution disease profiles while maintaining forecast stability and accuracy.
Mantis as a Foundation Model: Generalizing to New Mechanisms and Inputs
We next evaluated Mantis’s out-of-the-box forecasting capacity across both held-out transmission mechanisms and new input data modalities.
First, we benchmarked Mantis on diseases governed by transmission and progression mechanisms never represented in training simulations. For example, hepatitis B exhibits bloodborne transmission and chronic infection dynamics that are markedly different from the acute, self-limiting outbreaks used in pretraining. The Mantis training set included no models without recovery—all simulated diseases eventually transition out of the infectious state—and the longest average infectious period is just 10 days. In contrast, hepatitis B infection can persist for years. Despite this fundamental mechanistic mismatch, Mantis achieved competitive performance on hepatitis B forecasts, ranking second in MAE (0.85 vs 0.82) and achieving the best WIS (0.84) among all models (Table 1). These results demonstrate that Mantis can generalize across fundamental mechanistic boundaries—not just to new pathogens, but to qualitatively different epidemic regimes, such as chronic persistence versus outbreak spread.
Next, we tested Mantis’s ability to generalize across data modalities. Although trained using cases, hospitalizations, and deaths, Mantis can incorporate new input types at inference time. We evaluated Mantis on syndromic surveillance data (percent of outpatient visits for influenza-like illness), a data type never represented in training simulations. Despite this distribution shift, Mantis achieved the lowest MAE (0.83) and WIS (0.88) on ILI forecasting (Table 1), demonstrating robust generalization to novel surveillance indicators. This capability is especially valuable in settings where conventional case counts may be missing or delayed, but syndromic indicators are available in near real-time.
Across all seven disease forecasting tasks, spanning respiratory, vector-borne, bloodborne, and syndromic targets, Mantis consistently ranked first or second on both point and probabilistic metrics (Table 1). This level of generalization, across diseases, mechanisms, and modalities, demonstrates Mantis’s capacity as a foundation model and the advantages of scaling afforded by simulation data. Rather than relying on domain-specific tuning, Mantis captures structural features of contagion that transfer across contexts, enabling immediate deployment in settings where other models would require retraining, expert tuning, or disease-specific data collection.
Discussion
Mantis demonstrates that accurate forecasting across diverse diseases can be achieved through training on simulated outbreaks, without access to real-world data. By learning from mechanistic simulations, Mantis delivers competitive performance across real-world forecasting tasks spanning multiple transmission modes, outbreak phases, and surveillance systems. These capabilities position Mantis as a practical tool for public health use, especially in low-resource or emerging outbreak settings where data may be limited or delayed. Because Mantis can generate forecasts with only a short initial observation window, it enables early situational awareness when conventional models require longer histories to achieve comparable accuracy. Mantis’s advantage in data-scarce settings stems from its ability to avoid overfitting to sparse initial data. Rather than extrapolating confidently from a handful of observations, Mantis produces appropriately hedged forecasts that reflect the true uncertainty across potential outbreak trajectories, enabling more reliable probabilistic planning even when historical context is minimal.
The implications extend beyond public health. Simulation-supervised learning allows mechanistic models to embed scientific knowledge into neural networks, enabling them to generalize beyond observed data [8, 9]. This paradigm offers a potential solution to a fundamental challenge in epidemiological modeling: real-world disease data are limited by privacy concerns, reporting inconsistencies, and the infrequency of major outbreaks. Simulations, by contrast, can generate unlimited training data spanning diverse scenarios. Whether simulation-based training ultimately proves superior to training on massive real-world datasets—if such data could be assembled—remains an open question. However, given current data availability constraints, the simulation-based approach offers a practical path forward for building general-purpose forecasting systems.
Mantis represents an initial attempt at a foundation model for disease forecasting, and several important questions remain unresolved. First, the optimal neural architecture for this task is unclear. While our hybrid CNN-transformer design performs well, alternative architectures may offer better performance or efficiency trade-offs. Second, the ideal balance between simulation diversity and computational cost requires further investigation. Our training set spans 400 million simulated days, generated continuously until model convergence without repeating data. Using synthetic data like this allows us to avoid the overfitting risks inherent in epoch-based training on fixed datasets. However, whether this scale is necessary or whether comparable performance could be achieved with fewer but more strategically designed simulations remains to be seen.
While Mantis shows broad generalization, several limitations merit consideration. Though we demonstrate generalization beyond the specific regimes directly simulated during training, the model’s performance depends on the diversity and realism of training simulations. Additionally, Mantis inherits any systematic biases present in the underlying mechanistic models used for training. If the simulation framework systematically misrepresents certain disease dynamics—for example, categorically overestimating intervention effects—these biases will propagate to the trained model. Future work should explore approaches to identify and fill critical simulation gaps, and develop methods to validate that training simulations adequately span the space of real-world disease dynamics.
Mantis illustrates a path forward for general-purpose disease forecasting systems that can be quickly deployed without disease-specific tuning. Its out-of-the-box generalization makes it especially well-suited for early outbreak response, rural regions, and low- and middle-income countries—settings where traditional models often fail due to sparse or delayed data. By pairing simulation-based priors with flexible neural architectures, this approach offers a scalable alternative to expert-tuned, disease-specific models. As a first demonstration of simulation-grounded foundation models in infectious disease forecasting, Mantis establishes both the promise and the open challenges of this paradigm for future decision support systems in global public health.
Methods
A complete description of the methods is given in the Appendix.
Training Mantis on Simulated Outbreaks
Mantis was trained entirely on synthetic outbreaks generated using mechanistic epidemiological models, following the framework developed in [8]. Simulations were based on compartmental disease models with stochastic transitions and heterogeneous parameterizations, using three primary transmission routes: human-to-human respiratory illness models, vectorborne disease models (primarily inspired by arbovirus dynamics), and environmental transmission models (inspired by cholera and other environmentally transmitted pathogens). These models varied in transmission dynamics, disease severity, and intervention strategies to create a broad and diverse training distribution (see Appendix A for details). Each simulation generated daily time series of cases, hospitalizations, and deaths under a wide range of epidemic scenarios. In total, the training dataset included over 400 million days of data spanning thousands of parameter combinations.
Neural Architecture and Training Procedure
Mantis uses a sequence-to-sequence neural network with convolutional and transformer-based components (Appendix B). The model takes an input context window of up to 112 weeks of observed time series and outputs probabilistic forecasts for multiple future targets (e.g., cases, hospitalizations, deaths). Forecasts are trained using a weighted quantile loss. We optimized the neural network using AdamW [20]. No real-world data are used during training.
Evaluation on Real-World Forecasting Tasks
We evaluated Mantis across six diverse real-world infectious disease forecasting scenarios to assess generalization across pathogens, transmission modes, geographic scales, and historical periods (Appendix C). Our evaluation comprised two primary tasks and four additional challenges representing varying degrees of distribution shift from the model’s training data. Together, these datasets encompass over 39 forecasting tasks and more than 150,000 individual forecast evaluations, enabling rigorous assessment of Mantis’s capabilities across diverse epidemiological conditions, spatial resolutions, and outcome types. All Mantis evaluations were conducted in a strict zero-shot setting, with no fine-tuning or retraining on any evaluation data.
Baseline Models
We compared Mantis against two classes of baseline models: operational forecasts from the CDC COVID-19 Forecast Hub and models we re-implemented for evaluation across all diseases. For the CDC COVID-19 Forecast Hub comparison, we used forecasts as originally submitted by modeling teams during the pandemic and, to the best of our ability, evaluated both them and Mantis using the data as it was available at the time of each forecast (i.e., without any backfill or cleaning). This retrospective evaluation means our dataset contains data quality issues (e.g., missing/underreported data) that were present in the original surveillance data and likely affected all models’ performance, including Mantis. These CDC models were retrained or refitted each week to incorporate newly available data, and modeling teams could manually adjust forecasts if their automated systems produced anomalous predictions. Mantis forecasts were not modified in any way. This represents the strongest possible real-world baseline, reflecting operational forecasting practice with expert oversight under realistic data conditions.
For baselines we implemented—including statistical models (exponential smoothing and SARIMA) and deep learning models (LSTM)—we likewise retrained or refitted each model at every forecast point as new data became available. This rolling retraining procedure provides the strongest possible baseline performance for each method. Full baseline model implementation details can be found in Appendix C.
To compute relative MAE and relative WIS, we normalized all forecast errors by the performance of a naïve baseline model. This baseline generates point forecasts by projecting the most recent observed value forward to all future time points. Uncertainty intervals are constructed using the procedure from the COVIDhub-baseline model [2]: we collect all first differences from the historical time series along with their negatives (i.e., $y_{t}-y_{t-1}$ and $(y_{t}-y_{t-1})$ for all past times $t$). We form a piecewise linear approximation to the empirical cumulative distribution function of these differences, then generate 100,000 Monte Carlo samples by independently drawing sequences of $h$ differences and adding them to the most recent observed value. The resulting distribution is truncated at zero (to prevent negative forecasts) and summarized using quantiles, with the median set equal to the last observed value to eliminate sampling noise. This baseline captures the natural variability in historical changes without making assumptions about future trends, providing a neutral reference point for evaluating forecast skill.
Author Contributions
CD led the study and developed the code and initial draft of the manuscript. CH and AS did the ablation studies and RM developed figure visualizations. CD, MCE, RM, and EM all contributed substantial edits and revisions for the final manuscript.
Appendix A Training Data
Human-to-Human Transmission Simulator
The human-to-human transmission simulator implements a stochastic compartmental model that captures the complex dynamics of respiratory pathogens spreading through direct contact. This simulator generates diverse outbreak scenarios by randomly sampling from wide parameter ranges, producing synthetic epidemics that span the spectrum from influenza-like seasonal patterns to COVID-like pandemic dynamics.
Model Structure
Compartmental Framework
The core model follows a flexible SEAIR structure (Susceptible, Exposed, Asymptomatic, Infectious, Recovered) with optional latent periods, asymptomatic transmission, and waning immunity. These structural features are included probabilistically during simulation generation, allowing Mantis to learn from a wide range of outbreak regimes, including both acute and chronic infections, and both simple and complex transmission pathways. For example, in some simulations, all exposed individuals transition directly to the infectious state (i.e., no latent period); in others, asymptomatic infections are excluded, reflecting diseases without subclinical spread.
When fully enabled, the population is divided into five compartments:
$\frac{dS}{dt}$ $=-\lambda S+\omega R+bN-dS$ (1)
$\frac{dE}{dt}$ $=\lambda S-\sigma E-dE$ (2)
$\frac{dA}{dt}$ $=p_{a}\sigma E-\gamma_{a}A-dA$ (3)
$\frac{dI}{dt}$ $=(1-p_{a})\sigma E-\gamma I-dI$ (4)
$\frac{dR}{dt}$ $=\gamma_{a}A+\gamma I-\omega R-dR$ (5)
where:
  • $S$: Susceptible individuals
  • $E$: Exposed (latent) individuals
  • AA: Asymptomatic infectious individuals
  • II: Symptomatic infectious individuals
  • $R$: Recovered/immune individuals
  • $N=S+E+A+I+R$: Total population
Force of Infection
The force of infection $\lambda$ incorporates contributions from both symptomatic and asymptomatic individuals:
$\lambda(t)=\beta(t)\cdot s(t)\cdot\left(\frac{I+\alpha A}{N}\right)\cdot m(t)$ (6)
where:
  • $\beta(t)$: Time-varying transmission rate (allows for multiple waves)
  • $s(t)$: Seasonal forcing factor
  • $\alpha$: Relative transmissibility of asymptomatic individuals
  • $m(t)$: Super-spreading multiplier
Model Parameters
The model includes the following epidemiological parameters:
  • $\sigma$: Progression rate from exposed to infectious ($1/\sigma$ = latent period)
  • $\gamma$: Recovery rate for symptomatic individuals ($1/\gamma$ = infectious period)
  • $\gamma_{a}$: Recovery rate for asymptomatic individuals
  • $\omega$: Waning immunity rate ($1/\omega$ = duration of immunity)
  • $p_{a}$: Probability of asymptomatic infection
  • $b$: Birth rate
  • $d$: Death rate
Stochastic Implementation
The continuous-time model is implemented as a discrete-time stochastic process using daily time steps. At each time step $t$, transitions between compartments are drawn from binomial distributions (equivalent to tau-leaping with large tau, but we found that it improved performance):
New exposures $\sim\text{Binomial}(S_{t},1-e^{-\lambda_{t}})$ (7)
$\text{E}\to\text{I/A transitions}$ $\sim\text{Binomial}(E_{t},1-e^{-\sigma})$ (8)
Recoveries from I $\sim\text{Binomial}(I_{t},1-e^{-\gamma})$ (9)
Recoveries from A $\sim\text{Binomial}(A_{t},1-e^{-\gamma_{a}})$ (10)
Waning immunity $\sim\text{Binomial}(R_{t},1-e^{-\omega})$ (11)
Among newly infectious individuals, the split between asymptomatic and symptomatic follows:
New asymptomatic $\sim\text{Binomial}(\text{New infectious},p_{a})$ (12)
Multi-Wave Dynamics
To capture realistic outbreak patterns with multiple waves, the transmission rate $\beta(t)$ is implemented as a piecewise constant function:
$\beta(t)=\beta_{i}\quad\text{for }t\in[t_{i-1},t_{i})$ (13)
where $t_{0},t_{1},\ldots,t_{k}$ are randomly chosen wave change points and $\beta_{1},\beta_{2},\ldots,\beta_{k}$ are independently sampled transmission rates. The number of wave changes is drawn from a discrete uniform distribution, and change points are sampled uniformly within the simulation period.
Seasonal Forcing
Seasonal transmission patterns are modeled using multiple harmonic components:
$s(t)=1+\sum_{j=1}^{n_{h}}A_{j}\cos\left(\frac{2\pi(t-\phi_{j}+\delta_{y})}{P_{j}}\right)+\epsilon_{t}$ (14)
where:
  • $n_{h}$: Number of harmonic components (1-4)
  • $A_{j}$: Amplitude of harmonic $j$
  • $P_{j}$: Period of harmonic $j$ (365.0, 182.5, or 91.25 days)
  • $\phi_{j}$: Phase offset for harmonic $j$
  • $\delta_{y}$: Annual peak jitter, sampled independently each year to slightly shift seasonal peaks and avoid perfectly periodic recurrence
  • $\epsilon_{t}$: Daily multiplicative noise $\sim\mathcal{N}(1,0.05^{2})$
Super-Spreading Events
Super-spreading is modeled through a stochastic multiplier applied to the force of infection. At each time step, a subset of infectious individuals becomes super-spreaders with probability $p_{ss}$. The transmission enhancement follows a Gamma distribution:
$m(t)=\left(1-\frac{n_{ss}}{n_{I}}\right)+\frac{n_{ss}}{n_{I}}\cdot M$ (15)
where $n_{ss}\sim\text{Binomial}(n_{I},p_{ss})$ and $M\sim\text{Gamma}(k_{ss},\theta_{ss})$.
Intervention Policies
The simulator includes an optional universal intervention policy that modulates transmission based on case thresholds. The policy operates through a contact reduction factor applied to $\beta(t)$:
$\beta_{eff}(t)=\beta(t)\cdot r_{contact}(t)$ (16)
where $r_{contact}(t)\in[0.2,1.0]$ represents the intervention strength. The policy activates when daily cases exceed an on-threshold and deactivates after cases remain below an off-threshold for a specified duration.
Endemic Dynamics
For simulations with endemic circulation, demographic processes are included:
Births $\sim\text{Poisson}(b\cdot N)$ (17)
Deaths $\sim\text{Binomial}(X,\min(d,1))\quad\forall X\in\{S,E,A,I,R\}$ (18)
Importations $\sim\text{Poisson}(\iota)$ (19)
New births enter the susceptible compartment, deaths occur proportionally across all compartments, and imported cases enter the exposed or infectious compartment.
Observation Model
True epidemic dynamics are converted to observed case counts through a comprehensive noise injection pipeline that simulates real-world surveillance artifacts:
Multiplicative Noise
Base case counts are perturbed with log-normal noise:
$\hat{C}_{t}=C_{t}\cdot\epsilon_{t},\quad\epsilon_{t}\sim\text{LogNormal}(0,\sigma_{mult}^{2})$ (20)
Underreporting
Reporting rates evolve logistically over time:
$r_{t}=r_{0}+(r_{\infty}-r_{0})\cdot\frac{1}{1+e^{-k(t-t_{mid})}}$ (21)
Reported cases follow: $\tilde{C}_{t}\sim\text{Binomial}(\hat{C}_{t},r_{t})$
Reporting Delays
Cases are distributed across future days using inverse-weighted delay distributions:
$P(\text{delay}=d)\propto\frac{\alpha}{d+1},\quad d\in[0,D_{max}]$ (22)
Day-of-Week Effects
Reporting varies by weekday with multiplicative factors:
$\tilde{C}_{t}^{\prime}=\tilde{C}_{t}\cdot w(\text{day-of-week}(t))$ (23)
where $w(\cdot)$ ranges from 0.4 (weekend) to 1.2 (Tuesday).
Hospitalization and Death Generation
Secondary outcomes (hospitalizations and deaths) are derived from symptomatic cases using time-varying probabilities and delay distributions. For each wave segment $i$, hospitalization and death probabilities ($p_{h,i}$ and $p_{d,i}$) are independently sampled.
Hospitalization Process
$H_{t}$ $=\sum_{\tau=0}^{t}\text{Binomial}(I_{\tau},p_{h,wave(\tau)})\cdot f_{h}(t-\tau)$ (24)
$f_{h}(d)$ $\sim\text{Gamma}(k_{h},\theta_{h})$ (25)
Death Process
Deaths occur among hospitalized individuals with additional delays:
$D_{t}$ $=\sum_{\tau=0}^{t}\sum_{s=0}^{\tau}\text{Binomial}(H_{\tau,s},p_{d,wave(s)})\cdot f_{d}(t-\tau)$ (26)
$f_{d}(d)$ $\sim\text{Gamma}(k_{d},\theta_{d})$ (27)
where $H_{\tau,s}$ represents hospitalizations on day $\tau$ originating from infections on day $s$.
Parameter Ranges
Table LABEL:tab:h2h_params summarizes all parameters and their sampling distributions used in the human-to-human transmission simulator. Parameters were chosen to be broadly consistent with respirator pathogens.
Table 2: Human-to-Human Transmission Simulator Parameters
Parameter Symbol Range/Distribution Units
Population and Simulation
Population size NN Log-uniform[50,000, 40,000,000] individuals
Simulation days $T$ 2,000 days
Core Transmission Parameters
Base transmission rate $\beta$ Uniform[0.2, 0.235] day-1
Recovery rate $\gamma$ Uniform[0.1, 0.33] day-1
Latent progression rate $\sigma$ Uniform[0.2, 0.4] [21] day-1
Waning immunity rate $\omega$ Uniform[0.001, 0.0075] [22] day-1
Asymptomatic Dynamics
Asymptomatic probability $p_{a}$ Beta(3, 7) proportion
Asymp. transmissibility $\alpha$ Beta(2, 5) proportion
Has asymptomatic compartment - Bernoulli(0.5) boolean
Model Structure Options
Has latent period - Bernoulli(0.7) boolean
Has waning immunity - Bernoulli(0.9) boolean
Multi-Wave Dynamics
Multiple waves enabled - Configurable (default: True) boolean
Number of wave changes - Uniform[0, 4] count
Wave change days - Uniform[50, 1800] days
Super-Spreading
Super-spreader probability $p_{ss}$ Uniform[0.0005, 0.02] proportion
Multiplier shape $k_{ss}$ 4.0 -
Multiplier scale $\theta_{ss}$ 1.5 -
Seasonality
Use seasonality - Bernoulli(0.8) boolean
Number of harmonics $n_{h}$ Uniform[1, 4] count
Harmonic amplitude $A_{j}$ Base $\times$ Uniform[0.3, 1.0] proportion
Base amplitude - Uniform[0.1, 0.5] proportion
Phase offset $\phi_{j}$ Uniform[0, 365] + Uniform[-60, 60] days
Possible periods $P_{j}$ Choice[365.0, 182.5, 91.25] days
Daily noise std $\sigma_{\epsilon}$ 0.05 proportion
Annual peak jitter $\delta_{y}$ Uniform[-30, 30] days
Endemic Demographics
Enable endemic - Bernoulli(0.8) boolean
Birth rate $b$ Uniform[0.00002, 0.00012] day-1
Death rate factor - Uniform[0.8, 1.3] proportion
Importation rate $\iota$ Exp(Uniform[ln(0.01), ln(0.5)]) day-1
Intervention Policy
Enable intervention - Bernoulli(0.25) boolean
On threshold - $N\times$ Uniform[1e-5, 1e-3] cases/day
Off threshold - On threshold $\times$ Uniform[0, 1] cases/day
Contact reduction $r_{contact}$ Uniform[0.2, 0.6] proportion
Trigger delay - Uniform[0, 21] days
Minimum duration - Uniform[14, 35] days
Maximum duration - Choice[None, Uniform[60, 120]] days
Consecutive off days - Uniform[1, 50] days
Observation Noise
Multiplicative noise std $\sigma_{mult}$ 0.1 log-scale
Overdispersion parameter $r$ 1200 -
Reporting Effects
Apply reporting rate - Bernoulli(0.8) boolean
Initial reporting rate $r_{0}$ Uniform[0.05, 0.4] proportion
Final reporting rate $r_{\infty}$ Uniform[0.25, 0.85] proportion
Days to max testing - Uniform[30, 365] days
Logistic steepness $k$ Uniform[4, 8] -
Apply weekday effects - Bernoulli(0.8) boolean
Monday factor - Normal(1.2, 0.15) proportion
Tue-Thu factor - Normal(1.0, 0.1) proportion
Friday factor - Normal(0.9, 0.12) proportion
Saturday factor - Normal(0.6, 0.2) proportion
Sunday factor - Normal(0.4, 0.2) proportion
Apply lab noise - Bernoulli(0.8) boolean
Lab batch size - Poisson(100) samples
Bad batch rate - 0.005 proportion
Bad batch accuracy - Uniform[0.7, 0.85] proportion
Apply delays - Bernoulli(0.8) boolean
Initial max delay - Uniform[7, 21] days
Final max delay - Uniform[2, 7] days
Initial alpha $\alpha_{0}$ 1.0 -
Final alpha $\alpha_{\infty}$ 4.0 -
Hospitalization & Death
Hospitalization probability $p_{h,i}$ Uniform[0.02, 0.15] per wave proportion
Death probability (of hosp.) $p_{d,i}$ Uniform[0.05, 0.3] per wave proportion
Hosp. delay shape $k_{h}$ Uniform[2.0, 4.0] -
Hosp. delay scale $\theta_{h}$ Uniform[1.0, 3.0] days
Death delay shape $k_{d}$ Uniform[1.5, 2.5] -
Death delay scale $\theta_{d}$ Uniform[4.5, 8.5] days
This comprehensive parameter space enables the generation of diverse outbreak scenarios spanning different pathogen characteristics, population contexts, surveillance qualities, and intervention regimes. Each simulation run samples independently from these distributions, creating a rich training corpus that captures the full spectrum of human-to-human transmission dynamics.
Vector-Borne Transmission Simulator
The vector-borne transmission simulator implements a coupled host-vector compartmental model that captures the complex dynamics of diseases transmitted through arthropod vectors, such as dengue, Zika, chikungunya, and malaria. This simulator models the epidemiological dynamics within both human and vector populations, incorporating realistic vector ecology, seasonal patterns, and intervention strategies.
Model Structure
Coupled Host-Vector Framework
The model consists of two interacting population systems: humans and vectors (typically mosquitoes). The human population follows a SEAIR structure similar to the human-to-human model, while vectors follow a simpler SEI structure without recovery (as infected vectors typically remain infectious for life).
Human Population Dynamics
The human compartments evolve according to:
$\frac{dS_{h}}{dt}$ $=-\lambda_{h}S_{h}+\omega R_{h}+b_{h}N_{h}-d_{h}S_{h}$ (28)
$\frac{dE_{h}}{dt}$ $=\lambda_{h}S_{h}-\sigma_{h}E_{h}-d_{h}E_{h}$ (29)
$\frac{dA_{h}}{dt}$ $=p_{a}\sigma_{h}E_{h}-\gamma_{a}A_{h}-d_{h}A_{h}$ (30)
$\frac{dI_{h}}{dt}$ $=(1-p_{a})\sigma_{h}E_{h}-\gamma I_{h}-d_{h}I_{h}$ (31)
$\frac{dR_{h}}{dt}$ $=\gamma_{a}A_{h}+\gamma I_{h}-\omega R_{h}-d_{h}R_{h}$ (32)
Vector Population Dynamics
The vector compartments evolve according to:
$\frac{dS_{v}}{dt}$ $=\mu_{v}N_{v}-\lambda_{v}S_{v}-\mu_{v}S_{v}$ (33)
$\frac{dE_{v}}{dt}$ $=\lambda_{v}S_{v}-\sigma_{v}E_{v}-\mu_{v}E_{v}$ (34)
$\frac{dI_{v}}{dt}$ $=\sigma_{v}E_{v}-\mu_{v}I_{v}$ (35)
where $N_{v}=S_{v}+E_{v}+I_{v}$ is the total vector population.
Force of Infection
The transmission between humans and vectors occurs through biting interactions. The force of infection for humans (from infected vectors) is:
$\lambda_{h}(t)=a\cdot b_{h}\cdot\frac{I_{v}}{N_{v}}\cdot\beta(t)\cdot s(t)\cdot f_{intervention}(t)$ (36)
The force of infection for vectors (from infected humans) is:
$\lambda_{v}(t)=a\cdot b_{v}\cdot\frac{I_{h}+\alpha A_{h}}{N_{h}}\cdot\beta(t)\cdot s(t)\cdot f_{intervention}(t)$ (37)
where:
  • $a$: Biting rate (bites per vector per day)
  • $b_{h}$: Probability of transmission from vector to human per bite
  • $b_{v}$: Probability of transmission from human to vector per bite
  • $\beta(t)$: Time-varying transmission scaling factor
  • $s(t)$: Seasonal forcing function
  • $f_{intervention}(t)$: Intervention effectiveness factor
  • $\alpha$: Relative transmissibility of asymptomatic humans
Stochastic Implementation
The continuous-time model is implemented as a discrete-time stochastic process. At each time step $t$, transitions are drawn from binomial distributions:
Human Transitions
New human exposures $\sim\text{Binomial}(S_{h}(t),1-e^{-\lambda_{h}(t)})$ (38)
$\text{E}_{h}\to\text{I}_{h}/\text{A}_{h}$ $\sim\text{Binomial}(E_{h}(t),1-e^{-\sigma_{h}})$ (39)
Human recoveries from I $\sim\text{Binomial}(I_{h}(t),1-e^{-\gamma})$ (40)
Human recoveries from A $\sim\text{Binomial}(A_{h}(t),1-e^{-\gamma_{a}})$ (41)
Waning immunity $\sim\text{Binomial}(R_{h}(t),1-e^{-\omega})$ (42)
Vector Transitions
New vector exposures $\sim\text{Binomial}(S_{v}(t),1-e^{-\lambda_{v}(t)})$ (43)
$\text{E}_{v}\to\text{I}_{v}$ $\sim\text{Binomial}(E_{v}(t),1-e^{-\sigma_{v}})$ (44)
Vector deaths from S $\sim\text{Binomial}(S_{v}(t),\mu_{v})$ (45)
Vector deaths from E $\sim\text{Binomial}(E_{v}(t),\mu_{v})$ (46)
Vector deaths from I $\sim\text{Binomial}(I_{v}(t),\mu_{v})$ (47)
New vector births $\sim\text{Poisson}(\mu_{v}N_{v}(t))$ (48)
Vector Population Dynamics
Vector populations exhibit distinct ecological patterns compared to human populations:
Mortality and Reproduction
Vectors have high mortality rates ($\mu_{v}$) with typical lifespans of 10-30 days. To maintain population equilibrium, new vectors are born at rate $\mu_{v}N_{v}$, entering the susceptible class.
Seasonal Patterns
Vector abundance and activity are strongly seasonal, captured through enhanced seasonal forcing:
$s(t)=1.2+\sum_{j=1}^{n_{h}}A_{j}\cos\left(\frac{2\pi(t-\phi_{j}+\delta_{y})}{P_{j}}\right)+\epsilon_{t}$ (49)
The baseline value (1.2) and amplitude ranges ($A_{j}\in[0.5,1.25]$) are increased compared to human-to-human transmission to reflect the stronger seasonal dependence of vector-borne diseases.
Multi-Wave Dynamics
Similar to the human-to-human simulator, vector-borne outbreaks can exhibit multiple waves through time-varying transmission scaling:
$\beta(t)=\beta_{i}\quad\text{for }t\in[t_{i-1},t_{i})$ (50)
However, for vector-borne diseases, wave changes often reflect environmental conditions affecting vector populations (rainfall, temperature) or intervention campaigns rather than behavioral changes.
Intervention Modeling
Vector control interventions are modeled through the reduction factor $f_{intervention}(t)$ applied to both $\lambda_{h}$ and $\lambda_{v}$. Common interventions include:
  • Vector control: Reduces vector population or biting rates
  • Larvicide applications: Affects vector reproduction
  • Personal protection: Reduces effective contact rates
  • Environmental management: Modifies vector habitat suitability
The intervention policy triggers based on human case thresholds but affects the vector reduction parameter:
$f_{intervention}(t)=\begin{cases}f_{vector}&\text{if intervention active}\\ 1.0&\text{otherwise}\end{cases}$ (51)
where $f_{vector}\in[0.2,0.6]$ represents the effectiveness of vector control measures.
Initialization Strategy
Human Population
Initial conditions allow for pre-existing immunity in the human population:
Immune fraction $\sim\text{Uniform}(0,0.7)$ (52)
Initial infections $\sim\text{NegativeBinomial}(n=\max(1,N_{h}\times 5\times 10^{-6}),p=0.5)$ (53)
Vector Population
Vector populations start with a small fraction exposed/infected:
$N_{v}$ $=N_{h}\times\text{Uniform}(2.0,10.0)$ (54)
Initial vector exposures $\sim\text{NegativeBinomial}(n=\max(1,N_{v}\times 10^{-4}),p=0.3)$ (55)
The higher initial prevalence in vectors reflects their role as reservoirs and their shorter generation times.
Parameter Ranges
Table LABEL:tab:vector_params summarizes all parameters specific to the vector-borne transmission simulator.
Table 3: Vector-Borne Transmission Simulator Parameters
Parameter Symbol Range/Distribution Units
Human Population (Shared with H2H)
Population size $N_{h}$ Log-uniform[50,000, 40,000,000] individuals
Base transmission scaling $\beta$ Uniform[0.5, 0.7] -
Recovery rate $\gamma$ Uniform[0.1, 0.33] day-1
Latent progression rate $\sigma_{h}$ Uniform[0.2, 0.4] [21] day-1
Waning immunity rate $\omega$ Uniform[0.003, 0.02] [22] day-1
Asymptomatic probability $p_{a}$ Beta(3, 7) proportion
Asymp. transmissibility $\alpha$ Beta(2, 5) proportion
Vector Population Dynamics
Vector population size $N_{v}$ $N_{h}\times$ Uniform[2.0, 10.0] individuals
Vector mortality rate $\mu_{v}$ Uniform[0.03, 0.10] day-1
Vector latent progression $\sigma_{v}$ Uniform[0.15, 0.30] day-1
Vector lifespan $1/\mu_{v}$ Uniform[10, 33] days
Extrinsic incubation $1/\sigma_{v}$ Uniform[3.3, 6.7] days
Transmission Parameters
Biting rate $a$ Uniform[0.40, 0.80] bites/vector/day
Vector→human transmission $b_{h}$ Uniform[0.35, 0.75] probability
Human→vector transmission $b_{v}$ Uniform[0.35, 0.75] probability
Model Structure Options
Has latent period - Bernoulli(0.7) boolean
Has asymptomatic compartment - Bernoulli(0.5) boolean
Has waning immunity - Bernoulli(1.0) boolean
Enhanced Seasonality
Use seasonality - Bernoulli(0.8) boolean
Base seasonal value - 1.2 -
Harmonic amplitude range $A_{j}$ Base $\times$ Uniform[0.3, 1.0] proportion
Base amplitude - Uniform[0.5, 1.25] proportion
Number of harmonics $n_{h}$ Uniform[1, 4] count
Multi-Wave Dynamics
Multiple waves enabled - Configurable (default: True) boolean
Number of wave changes - Uniform[0, 4] count
Wave change timing - Uniform[50, 1800] days
Vector Control Interventions
Enable intervention - Bernoulli(0.25) boolean
On threshold - $N_{h}\times$ Uniform[1e-5, 1e-3] cases/day
Off threshold - On threshold $\times$ Uniform[0, 1] cases/day
Vector reduction $f_{vector}$ Uniform[0.2, 0.6] proportion
Trigger delay - Uniform[0, 21] days
Minimum duration - Uniform[14, 35] days
Maximum duration - Choice[None, Uniform[60, 120]] days
Consecutive off days - Uniform[1, 50] days
Initialization Parameters
Human immune fraction - Uniform[0.0, 0.7] proportion
[…]
Vector:human ratio - Uniform[2.0, 10.0] ratio
[…]
Observation & Outcomes (Shared)
Hospitalization probability $p_{h}$ Uniform[0.05, 0.20] proportion
Death probability (of hosp.) $p_{d}$ Uniform[0.10, 0.40] proportion
Hosp. delay shape $k_{h}$ Uniform[2.0, 4.0] -
Hosp. delay scale $\theta_{h}$ Uniform[1.0, 3.0] days
Death delay shape $k_{d}$ Uniform[1.5, 2.5] -
Death delay scale $\theta_{d}$ Uniform[2.0, 5.0] days
Noise & Reporting (Identical to H2H)
Multiplicative noise std $\sigma_{mult}$ 0.1 log-scale
Reporting effects - Same as Table LABEL:tab:h2h_params various
Day-of-week effects - Same as Table LABEL:tab:h2h_params various
Lab noise parameters - Same as Table LABEL:tab:h2h_params various
Delay distributions - Same as Table LABEL:tab:h2h_params various
Key Differences from Human-to-Human Transmission
The vector-borne simulator incorporates several critical differences that distinguish it from direct human-to-human transmission:
Dual Population Dynamics
Unlike human-to-human models, vector-borne diseases require explicit modeling of vector population dynamics, including birth, death, and infection processes within the arthropod population.
Transmission Pathway
Transmission requires a two-step process (human → vector → human) rather than direct contact, fundamentally altering the epidemic dynamics and introducing additional delay and complexity.
Environmental Sensitivity
Vector-borne diseases exhibit stronger seasonal patterns due to vector ecology dependence on temperature, rainfall, and other environmental factors, reflected in enhanced seasonal forcing parameters.
Vector-Specific Interventions
Control strategies focus on vector population reduction rather than behavioral modifications, affecting both $\lambda_{h}$ and $\lambda_{v}$ through environmental rather than social mechanisms.
Basic Reproduction Number
$R_{0}$ calculations must account for the round-trip transmission cycle and vector population dynamics, resulting in a more complex formulation that depends on vector:human ratios and vector survival probabilities.
This comprehensive parameter space enables realistic simulation of diverse vector-borne disease scenarios, capturing the full spectrum of arthropod-transmitted pathogen dynamics across different ecological contexts, intervention strategies, and surveillance systems.
Environmental Transmission Simulator
The environmental transmission simulator implements a hybrid compartmental model that captures diseases transmitted through both direct human-to-human contact and indirect environmental pathways via contaminated water sources. This simulator models pathogens such as cholera, typhoid, hepatitis A, and norovirus, which exhibit complex transmission dynamics involving multiple routes of exposure and environmental persistence.
Model Structure
Dual-Route Transmission Framework
The waterborne model extends the basic SEAIR structure to include an additional environmental compartment representing pathogen concentration in water. The model simultaneously tracks both direct contact transmission (similar to respiratory diseases) and indirect transmission through contaminated water or environmental sources.
Human Population Dynamics
The human compartments follow the same structure as the human-to-human model:
$\frac{dS}{dt}$ $=-\lambda_{h}S+\omega R+bN-dS$ (56)
$\frac{dE}{dt}$ $=\lambda_{h}S-\sigma E-dE$ (57)
$\frac{dA}{dt}$ $=p_{a}\sigma E-\gamma_{a}A-dA$ (58)
$\frac{dI}{dt}$ $=(1-p_{a})\sigma E-\gamma I-dI$ (59)
$\frac{dR}{dt}$ $=\gamma_{a}A+\gamma I-\omega R-dR$ (60)
Environmental Pathogen Dynamics
The environmental water compartment evolves according to:
$\frac{dW}{dt}$ $=\eta_{I}I+\eta_{A}A-\mu_{w}W$ (61)
where:
  • $W$: Pathogen concentration in water sources
  • $\eta_{I}$: Contamination rate from symptomatic individuals
  • $\eta_{A}$: Contamination rate from asymptomatic individuals
  • $\mu_{w}$: Pathogen decay/removal rate in water
Composite Force of Infection
The total force of infection combines both transmission routes:
$\lambda_{h}(t)=\lambda_{contact}(t)+\lambda_{water}(t)$ (62)
where:
$\lambda_{contact}(t)$ $=\beta(t)\cdot s(t)\cdot\frac{I+\alpha A}{N}\cdot f_{contact}(t)$ (63)
$\lambda_{water}(t)$ $=\delta(t)\cdot s(t)\cdot W\cdot f_{water}(t)$ (64)
The parameters are:
  • $\beta(t)$: Time-varying contact transmission rate
  • $\delta(t)$: Time-varying waterborne transmission rate
  • $s(t)$: Seasonal forcing factor (affects both routes)
  • $f_{contact}(t)$, $f_{water}(t)$: Intervention effectiveness factors
  • $\alpha$: Relative transmissibility of asymptomatic individuals
Multi-Route Wave Dynamics
Waterborne diseases can exhibit complex temporal patterns due to independent variation in both transmission routes. The model implements time-varying transmission coefficients for both pathways:
$\beta(t)$ $=\beta_{i}\quad\text{for }t\in[t_{i-1},t_{i})$ (65)
$\delta(t)$ $=\delta_{i}\quad\text{for }t\in[t_{i-1},t_{i})$ (66)
Wave changes can reflect:
  • Contact route changes: Behavioral modifications, crowding patterns, hygiene practices
  • Water route changes: Infrastructure failures, seasonal water quality, sanitation disruptions
  • Coordinated changes: Public health interventions affecting both routes simultaneously
Environmental Contamination Process
Contamination Sources
Both symptomatic and asymptomatic individuals contribute to environmental contamination:
$\text{Daily contamination}=\eta_{I}\cdot I+\eta_{A}\cdot A\cdot\alpha_{env}$ (67)
where $\eta_{I}$ and $\eta_{A}$ represent base contamination rates, and $\alpha_{env}$ is the relative environmental shedding from asymptomatic individuals.
Pathogen Persistence and Decay
Environmental pathogens undergo exponential decay:
$\frac{dW}{dt}\Big|_{\text{decay}}=-\mu_{w}W$ (68)
The decay rate $\mu_{w}$ encompasses multiple removal mechanisms:
  • Natural pathogen die-off
  • Water treatment processes
  • Dilution effects
  • UV radiation and temperature effects
Stochastic Implementation
The continuous-time model is implemented stochastically with daily time steps:
Human Transmission Events
Contact exposures $\sim\text{Binomial}(S(t),1-e^{-\lambda_{contact}(t)})$ (69)
Water exposures $\sim\text{Binomial}(S(t),1-e^{-\lambda_{water}(t)})$ (70)
Total new exposures $=\text{Contact exposures}+\text{Water exposures}$ (71)
Note: In the stochastic implementation, total exposures are calculated as:
Total exposures $\sim\text{Binomial}(S(t),1-e^{-(\lambda_{contact}(t)+\lambda_{water}(t))})$ (72)
Environmental Process Updates
New contamination $=\eta_{I}\cdot I(t)+\eta_{A}\cdot A(t)\cdot\alpha_{env}$ (73)
Pathogen decay $=\mu_{w}\cdot W(t)$ (74)
$W(t+1)$ $=\max(0,W(t)+\text{New contamination}-\text{Pathogen decay})$ (75)
Compartment Transitions
Human compartment transitions follow the same binomial processes as the human-to-human model:
$\text{E}\to\text{I/A}$ $\sim\text{Binomial}(E(t),1-e^{-\sigma})$ (76)
Recovery from I $\sim\text{Binomial}(I(t),1-e^{-\gamma})$ (77)
Recovery from A $\sim\text{Binomial}(A(t),1-e^{-\gamma_{a}})$ (78)
Waning immunity $\sim\text{Binomial}(R(t),1-e^{-\omega})$ (79)
Enhanced Seasonality
Waterborne diseases exhibit strong seasonal patterns due to environmental factors affecting both transmission routes and pathogen survival:
$s(t)=1.0+\sum_{j=1}^{n_{h}}A_{j}\cos\left(\frac{2\pi(t-\phi_{j}+\delta_{y})}{P_{j}}\right)+\epsilon_{t}$ (80)
Seasonal Factors
Seasonality in waterborne diseases reflects:
  • Rainfall patterns: Affecting water contamination and treatment efficacy
  • Temperature effects: Influencing pathogen survival in water
  • Flooding events: Disrupting sanitation infrastructure
  • Social patterns: Seasonal migration, gatherings, and water usage
Enhanced Amplitude Range
Waterborne diseases show stronger seasonal variation with amplitude parameters:
$A_{j}\in[0.6,1.5]\text{ (compared to }[0.1,0.5]\text{ for human-to-human)}$ (81)
Dual-Route Intervention Modeling
Interventions can target either or both transmission routes with different effectiveness:
$\lambda_{contact,eff}(t)$ $=\lambda_{contact}(t)\cdot f_{contact}(t)$ (82)
$\lambda_{water,eff}(t)$ $=\lambda_{water}(t)\cdot f_{water}(t)$ (83)
Contact Route Interventions
Similar to human-to-human diseases:
  • Social distancing measures
  • Hygiene promotion
  • Isolation of cases
Water Route Interventions
Specific to environmental transmission:
  • Water treatment and chlorination
  • Sanitation infrastructure improvements
  • Safe water distribution
  • Waste management systems
Intervention Policy
The universal intervention policy can independently affect both routes:
$f_{contact}(t)$ $\in[0.2,0.6]\text{ when intervention active}$ (84)
$f_{water}(t)$ $\in[0.3,0.7]\text{ when intervention active}$ (85)
Super-Spreading Events
Super-spreading affects only the contact transmission route, as waterborne transmission is mediated by environmental persistence rather than individual variation:
$\lambda_{contact,ss}(t)=\lambda_{contact}(t)\cdot m_{ss}(t)$ (86)
where $m_{ss}(t)$ follows the same gamma-distributed multiplier as in human-to-human transmission.
Initialization Strategy
Human Population
Initial conditions allow for higher pre-existing immunity due to endemic circulation:
Immune fraction $\sim\text{Beta}(10,3.5)\text{ (higher mean than H2H)}$ (87)
Initial infections $\sim\text{NegativeBinomial}(n=N_{h}\times 10^{-6},p=0.5)$ (88)
Environmental Pathogen Load
The water compartment initializes at zero, building up through contamination from initial infections:
$W(0)=0$ (89)
Parameter Ranges
Table LABEL:tab:waterborne_params summarizes all parameters specific to the waterborne transmission simulator.
Table 4: Waterborne Transmission Simulator Parameters
Parameter Symbol Range/Distribution Units
Human Population (Shared)
Population size NN Log-uniform[50,000, 40,000,000] individuals
Recovery rate $\gamma$ Uniform[0.1, 0.33] day-1
Latent progression rate $\sigma$ Uniform[0.2, 0.4] day-1
Waning immunity rate $\omega$ Uniform[0.001, 0.01] day-1
Asymptomatic probability $p_{a}$ Beta(3, 7) proportion
Asymp. transmissibility $\alpha$ Beta(2, 5) proportion
Dual-Route Transmission
Contact transmission rate $\beta$ Uniform[0.0, 0.05] day-1
Water transmission rate $\delta$ Uniform[0.0005, 0.01] day-1
Water contamination rate $\eta$ Uniform[0.001, 0.01] day-1
Pathogen decay rate $\mu_{w}$ Uniform[0.05, 0.3] day-1
Pathogen half-life $\ln(2)/\mu_{w}$ Uniform[2.3, 13.9] days
Model Structure Options
Has latent period - Bernoulli(0.7) boolean
Has asymptomatic compartment - Bernoulli(0.6) boolean
Has waning immunity - Bernoulli(0.9) boolean
Multi-Wave Dynamics
Multiple waves enabled - Configurable (default: True) boolean
Number of wave changes - Uniform[0, 7] count
Wave change timing - Uniform[50, 1800] days
Contact wave segments - Wave count + 1 count
Water wave segments - Wave count + 1 count
Enhanced Seasonality
Use seasonality - Bernoulli(0.9) boolean
Number of harmonics $n_{h}$ Uniform[1, 4] count
Base amplitude - Uniform[0.6, 1.5] proportion
Harmonic amplitude $A_{j}$ Base $\times$ Uniform[0.3, 1.0] proportion
Phase offset $\phi_{j}$ Uniform[0, 365] + Uniform[-60, 60] days
Possible periods $P_{j}$ Choice[365.0, 182.5, 91.25] days
Daily noise std $\sigma_{\epsilon}$ 0.05 proportion
Annual peak jitter $\delta_{y}$ Uniform[-30, 30] days
Dual-Route Interventions
Enable intervention - Bernoulli(0.25) boolean
On threshold - $N\times$ Uniform[1e-5, 1e-3] cases/day
Off threshold - On threshold $\times$ Uniform[0, 1] cases/day
Contact reduction $f_{contact}$ Uniform[0.2, 0.6] proportion
Water reduction $f_{water}$ Uniform[0.3, 0.7] proportion
Trigger delay - Uniform[0, 21] days
Minimum duration - Uniform[14, 35] days
Maximum duration - Choice[None, Uniform[60, 120]] days
Consecutive off days - Uniform[1, 50] days
Super-Spreading (Contact Only)
Super-spreader probability $p_{ss}$ Uniform[0.0002, 0.02] proportion
Multiplier shape $k_{ss}$ 4.0 -
Multiplier scale $\theta_{ss}$ 1.5 -
Endemic Demographics
Enable endemic - Bernoulli(0.8) boolean
Birth rate $b$ Uniform[0.00002, 0.00012] day-1
Death rate factor - Uniform[0.8, 1.3] proportion
Importation rate $\iota$ Exp(Uniform[ln(0.01), ln(0.5)]) day-1
Initialization Parameters
Novel outbreak probability - 0.0 (always endemic) proportion
Immune fraction (if endemic) - Beta(10, 3.5) proportion
[…]
Initial water pathogen load $W(0)$ 0.0 concentration
Observation & Outcomes
Hospitalization probability $p_{h}$ Uniform[0.02, 0.15] proportion
Death probability (of hosp.) $p_{d}$ Uniform[0.05, 0.3] proportion
Hosp. delay shape $k_{h}$ Uniform[2.0, 4.0] -
Hosp. delay scale $\theta_{h}$ Uniform[1.0, 3.0] days
Death delay shape $k_{d}$ Uniform[1.5, 2.5] -
Death delay scale $\theta_{d}$ Uniform[2.0, 5.0] days
Noise & Reporting (Identical to H2H)
Multiplicative noise std $\sigma_{mult}$ 0.1 log-scale
Overdispersion parameter $r$ 100 -
Reporting effects - Same as Table LABEL:tab:h2h_params various
Day-of-week effects - Same as Table LABEL:tab:h2h_params various
Lab noise parameters - Same as Table LABEL:tab:h2h_params various
Delay distributions - Same as Table LABEL:tab:h2h_params various
Key Differences from Other Transmission Modes
The waterborne simulator incorporates several distinguishing features:
Dual Transmission Pathways
Unlike single-route models, waterborne diseases require simultaneous modeling of direct contact and environmental transmission, each with independent temporal dynamics and intervention susceptibilities.
Environmental Persistence
The explicit water compartment captures pathogen persistence in the environment, creating temporal delays and amplification effects not present in direct transmission models.
Route-Specific Interventions
Control strategies must address both transmission pathways, with water-focused interventions (treatment, sanitation) having different effectiveness profiles than contact-focused measures.
Enhanced Seasonal Dependence
Environmental factors create stronger seasonal patterns than purely social transmission modes, reflected in increased amplitude ranges and environmental sensitivity.
Endemic Circulation Patterns
Higher baseline immunity levels reflect the endemic nature of many waterborne diseases in affected populations, influencing outbreak dynamics and intervention effectiveness.
This comprehensive dual-route framework enables realistic simulation of waterborne disease dynamics across diverse environmental contexts, intervention strategies, and population immunity profiles, capturing the complex interplay between direct and indirect transmission pathways characteristic of environmentally-mediated pathogens.
Appendix B Model Architecture
Mantis is implemented as a hybrid convolutional-transformer architecture specifically designed for epidemiological time series forecasting. The model consists of five main components: (1) multi-scale input embeddings, (2) hybrid CNN-transformer encoder blocks, (3) an epidemic pattern memory bank, (4) an autoregressive quantile decoder, and (5) comprehensive temporal feature engineering. This section provides detailed specifications for each component.
B.1 Input Processing and Feature Engineering
B.1.1 Multi-Scale Convolutional Embedding
The input embedding layer processes time series data using parallel convolutions at multiple temporal scales to capture epidemiological patterns across different time horizons. For input tensor $\mathbf{X}\in\mathbb{R}^{B\times L\times 2}$ where $B$ is batch size, $L$ is sequence length, and the feature dimension includes both target values and covariates (e.g., deaths and hospitalizations), the multi-scale embedding applies four parallel 1D convolutions:
$\mathbf{H}_{\text{short}}$ $=\text{Conv1D}(\mathbf{X};k=3,d_{\text{out}}=d/4)$ (90)
$\mathbf{H}_{\text{med}}$ $=\text{Conv1D}(\mathbf{X};k=7,d_{\text{out}}=d/4)$ (91)
$\mathbf{H}_{\text{long}}$ $=\text{Conv1D}(\mathbf{X};k=15,d_{\text{out}}=d/4)$ (92)
$\mathbf{H}_{\text{vlong}}$ $=\text{Conv1D}(\mathbf{X};k=31,d_{\text{out}}=d/4)$ (93)
where $k$ denotes kernel size, $d$ is the target hidden dimension (default 512), and each convolution uses appropriate padding to maintain sequence length. Each scale captures different temporal patterns:
  • Short-term (k=3): Weekly variations and immediate trends
  • Medium-term (k=7): Monthly patterns and intervention responses
  • Long-term (k=15): Seasonal cycles and extended epidemic phases
  • Very long-term (k=31): Annual seasonality and yearly trends
Each convolutional output is independently normalized and activated:
$\tilde{\mathbf{H}}_{i}$ $=\text{Dropout}(\text{SiLU}(\text{LayerNorm}(\mathbf{H}_{i})))\quad\text{for }i\in\{\text{short, med, long, vlong}\}$ (94)
The multi-scale features are concatenated and projected to the target hidden dimension:
$\mathbf{H}_{\text{conv}}$ $=\text{LayerNorm}(\mathbf{W}_{\text{proj}}\cdot\text{Concat}([\tilde{\mathbf{H}}_{\text{short}},\tilde{\mathbf{H}}_{\text{med}},\tilde{\mathbf{H}}_{\text{long}},\tilde{\mathbf{H}}_{\text{vlong}}]))$ (95)
where $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{d\times d}$ is a learned projection matrix.
B.1.2 Static and Temporal Feature Embeddings
Mantis incorporates rich contextual information through multiple embedding layers:
Disease Type Embedding
Disease categories are embedded using a learned lookup table:
$\mathbf{e}_{\text{disease}}=\text{Embedding}(\text{disease\_type};|V|=3,d_{\text{disease}}=64)$ (96)
where the three disease types correspond to human-to-human, waterborne, and vector-borne transmission modes from the training simulations.
Population Embedding
Population size is processed through a multi-layer perceptron after log-normalization:
$\mathbf{e}_{\text{pop}}=\text{GELU}(\text{LayerNorm}(\mathbf{W}_{\text{pop}}\cdot\log(1+\text{population})))$ (97)
where $\mathbf{W}_{\text{pop}}\in\mathbb{R}^{64\times 1}$ projects the scalar population to a 64-dimensional embedding.
Enhanced Temporal Embeddings
Temporal patterns are captured through multiple cyclic embeddings derived from absolute day indices. For day index $d_{t}$ at timestep $t$:
$\mathbf{e}_{\text{dow}}$ $=\text{LayerNorm}(\text{Embedding}(d_{t}\bmod 7;|V|=7,d=128))$ (98)
$\mathbf{e}_{\text{month}}$ $=\text{LayerNorm}(\text{Embedding}(\lfloor d_{t}/30\rfloor\bmod 12;|V|=12,d=128))$ (99)
$\mathbf{e}_{\text{doy}}$ $=\text{LayerNorm}(\text{Embedding}(d_{t}\bmod 366;|V|=366,d=128))$ (100)
These embeddings capture day-of-week effects (reporting cycles), monthly seasonality, and annual patterns respectively.
B.1.3 Feature Integration
All embedding components are concatenated and projected to the model’s hidden dimension. Static embeddings are expanded across the temporal dimension:
$\mathbf{H}_{\text{input}}$ $=\text{LayerNorm}(\mathbf{W}_{\text{feat}}\cdot\text{Concat}([\mathbf{H}_{\text{conv}},\mathbf{E}_{\text{disease}},\mathbf{E}_{\text{pop}},\mathbf{e}_{\text{dow}},\mathbf{e}_{\text{month}},\mathbf{e}_{\text{doy}}]))$ (101)
where $\mathbf{E}_{\text{disease}},\mathbf{E}_{\text{pop}}\in\mathbb{R}^{B\times L\times d_{\text{embed}}}$ are the temporally expanded static embeddings, and $\mathbf{W}_{\text{feat}}\in\mathbb{R}^{d\times(d+64+64+128+128+128)}$ projects the concatenated features to hidden dimension $d$.
B.2 Hybrid CNN-Transformer Encoder
The encoder consists of $N$ stacked hybrid blocks (default $N=16$), each combining local convolutional processing with global attention mechanisms.
B.2.1 Temporal Pattern Attention
Each encoder block uses a custom attention mechanism (‘TemporalPatternAttention‘) that incorporates relative positional encoding specifically designed for temporal sequences:
$\mathbf{Q},\mathbf{K},\mathbf{V}$ $=\mathbf{H}\mathbf{W}_{Q},\mathbf{H}\mathbf{W}_{K},\mathbf{H}\mathbf{W}_{V}$ (102)
$\mathbf{A}_{i,j}$ $=\frac{\mathbf{Q}_{i}\mathbf{K}_{j}^{T}}{\sqrt{d_{k}}}+\mathbf{R}_{i-j+L-1}$ (103)
$\text{Attention}(\mathbf{H})$ $=\text{softmax}(\mathbf{A})\mathbf{V}$ (104)
where $\mathbf{R}\in\mathbb{R}^{2L-1\times h}$ is a learned relative positional bias matrix with $h$ attention heads, and $\mathbf{R}_{i-j+L-1}$ provides position-dependent bias for the attention score between positions $i$ and $j$.
The relative positional encoding allows the model to learn temporal relationships that are translation-invariant, which is crucial for epidemiological patterns that may occur at different absolute times but with similar relative timing.
B.2.2 CNN-Transformer Hybrid Block
Each encoder block (‘CNNTransformerBlock‘) follows a three-stage design with post-normalization:
Stage 1: Local Pattern Extraction
$\mathbf{H}^{(1)}$ $=\text{LayerNorm}(\mathbf{H}^{(0)}+\text{Dropout}(\text{LocalConv}(\mathbf{H}^{(0)})))$ (105)
where LocalConv applies depthwise separable convolution:
$\text{LocalConv}(\mathbf{H})$ $=\text{Conv1D}(\text{DepthwiseConv1D}(\mathbf{H};k=5);k=1)$ (106)
The depthwise convolution (kernel size 5) captures local temporal dependencies within each feature channel, followed by a pointwise convolution that mixes information across channels.
Stage 2: Global Context Attention
$\mathbf{H}^{(2)}$ $=\text{LayerNorm}(\mathbf{H}^{(1)}+\text{Dropout}(\text{TemporalPatternAttention}(\mathbf{H}^{(1)},\mathbf{M})))$ (107)
where $\mathbf{M}\in\{0,1\}^{B\times L}$ is the padding mask indicating valid sequence positions.
Stage 3: Position-wise Feed-Forward
$\mathbf{H}^{(3)}$ $=\text{LayerNorm}(\mathbf{H}^{(2)}+\text{Dropout}(\text{FFN}(\mathbf{H}^{(2)})))$ (108)
where the feed-forward network is:
$\text{FFN}(\mathbf{H})$ $=\mathbf{W}_{2}\text{GELU}(\mathbf{W}_{1}\mathbf{H}+\mathbf{b}_{1})+\mathbf{b}_{2}$ (109)
with $\mathbf{W}_{1}\in\mathbb{R}^{d\times d_{\text{ffn}}}$, $\mathbf{W}_{2}\in\mathbb{R}^{d_{\text{ffn}}\times d}$, and $d_{\text{ffn}}=2048$ by default.
B.3 Epidemic Pattern Memory Bank
The pattern memory module (‘EpidemicPatternMemory’) implements a learnable memory bank that stores prototypical epidemic patterns discovered during training.
B.3.1 Pattern Bank Initialization
The pattern bank consists of $K=256$ learnable pattern vectors:
$\mathbf{P}=[\mathbf{p}_{1},\mathbf{p}_{2},\ldots,\mathbf{p}_{K}]\in\mathbb{R}^{K\times d}$ (110)
Pattern vectors are initialized with small random values: $\mathbf{p}_{k}\sim\mathcal{N}(0,0.02^{2}\mathbf{I})$.
B.3.2 Pattern Matching and Retrieval
For encoder output $\mathbf{H}_{\text{enc}}\in\mathbb{R}^{B\times L\times d}$, pattern matching computes attention weights over the pattern bank:
$\mathbf{S}$ $=\text{FFN}_{\text{match}}(\text{LayerNorm}(\mathbf{H}_{\text{enc}}))\in\mathbb{R}^{B\times L\times K}$ (111)
$\mathbf{A}_{\text{pattern}}$ $=\text{softmax}(\mathbf{S}\odot\mathbf{M}_{\text{exp}})$ (112)
$\mathbf{H}_{\text{retrieved}}$ $=\mathbf{A}_{\text{pattern}}\mathbf{P}\in\mathbb{R}^{B\times L\times d}$ (113)
where $\mathbf{M}_{\text{exp}}\in\mathbb{R}^{B\times L\times 1}$ is the expanded padding mask, $\odot$ denotes element-wise multiplication, and $\text{FFN}_{\text{match}}$ is a two-layer network:
$\text{FFN}_{\text{match}}(\mathbf{x})=\mathbf{W}_{3}\text{GELU}(\mathbf{W}_{4}\mathbf{x})$ (114)
with $\mathbf{W}_{4}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{3}\in\mathbb{R}^{d\times K}$.
B.3.3 Pattern Integration
Retrieved patterns are projected and integrated with the encoder output using residual connections:
$\mathbf{H}_{\text{memory}}$ $=\text{LayerNorm}(\mathbf{H}_{\text{enc}}+\text{Dropout}(\mathbf{W}_{\text{proj}}\mathbf{H}_{\text{retrieved}}))$ (115)
where $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{d\times d}$ is a learned projection matrix.
B.4 Autoregressive Quantile Decoder
The decoder generates probabilistic forecasts using an autoregressive approach with a two-layer GRU and cross-attention to encoder memory.
B.4.1 Decoder Initialization
The decoder state is initialized using masked pooling of the encoder output:
$\mathbf{h}_{0}$ $=\mathbf{W}_{\text{init}}\cdot\text{MaskedPool}(\mathbf{H}_{\text{memory}},\mathbf{M})$ (116)
$\text{MaskedPool}(\mathbf{H},\mathbf{M})$ $=\frac{\sum_{t=1}^{L}\mathbf{M}_{:,t}\odot\mathbf{H}_{:,t,:}}{\sum_{t=1}^{L}\mathbf{M}_{:,t}+\epsilon}$ (117)
where $\mathbf{W}_{\text{init}}\in\mathbb{R}^{d\times d}$ and $\epsilon=10^{-10}$ prevents division by zero. The initial state is replicated for the two-layer GRU: $\mathbf{h}_{0}^{(1)}=\mathbf{h}_{0}^{(2)}=\mathbf{h}_{0}$.
B.4.2 Autoregressive Generation
For each forecast step $\tau\in\{1,2,\ldots,H\}$ where $H$ is the forecast horizon:
Step 1: Input Processing
The previous prediction $y_{\tau-1}$ is embedded:
$\mathbf{e}_{\tau}=\text{GELU}(\text{LayerNorm}(\mathbf{W}_{\text{input}}y_{\tau-1}))\in\mathbb{R}^{B\times d/4}$ (118)
Step 2: Cross-Attention
The top GRU layer output attends to encoder memory:
$\mathbf{q}_{\tau}$ $=(\mathbf{h}_{\tau-1}^{(2)})^{T}\in\mathbb{R}^{B\times 1\times d}$ (119)
$\mathbf{c}_{\tau}$ $=\text{MultiHeadAttention}(\mathbf{q}_{\tau},\mathbf{H}_{\text{memory}},\mathbf{H}_{\text{memory}},\neg\mathbf{M})$ (120)
where $\neg\mathbf{M}$ serves as the key padding mask (True for padded positions).
Step 3: GRU Update
The concatenated input is processed through the two-layer GRU:
$\mathbf{x}_{\tau}$ $=\text{Concat}([\mathbf{c}_{\tau},\mathbf{e}_{\tau}])\in\mathbb{R}^{B\times 1\times(d+d/4)}$ (121)
$\mathbf{h}_{\tau}^{(1)},\mathbf{h}_{\tau}^{(2)}$ $=\text{GRU}(\mathbf{x}_{\tau},[\mathbf{h}_{\tau-1}^{(1)},\mathbf{h}_{\tau-1}^{(2)}])$ (122)
Step 4: Quantile Prediction
23 quantiles are predicted using separate projection heads:
$\hat{q}_{\tau}^{(i)}$ $=\mathbf{W}_{\text{out}}^{(i)}\text{GELU}(\text{LayerNorm}(\mathbf{W}_{\text{hidden}}^{(i)}\mathbf{h}_{\tau}^{(2)}))$ (123)
$\quad\text{for }i\in\{1,2,\ldots,23\}$ (124)
where each quantile head has parameters $\mathbf{W}_{\text{hidden}}^{(i)}\in\mathbb{R}^{(d/2)\times d}$ and $\mathbf{W}_{\text{out}}^{(i)}\in\mathbb{R}^{1\times(d/2)}$.
The 23 quantiles are: [0.01, 0.025, 0.05, 0.1, 0.15, 0.2, 0.25, 0.3, 0.35, 0.4, 0.45, 0.5, 0.55, 0.6, 0.65, 0.7, 0.75, 0.8, 0.85, 0.9, 0.95, 0.975, 0.99].
B.4.3 Teacher Forcing
During training, teacher forcing is applied with linearly decaying probability:
$p_{\text{tf}}(s)$ $=p_{\text{initial}}\cdot\left(1-\frac{s}{S_{\text{total}}}\right)+p_{\text{final}}\cdot\frac{s}{S_{\text{total}}}$ (125)
where $s$ is the current training step, $S_{\text{total}}$ is the total number of training steps, $p_{\text{initial}}=0.2$, and $p_{\text{final}}=0.0$.
At each step, with probability $p_{\text{tf}}(s)$, the true target value is used as input instead of the model’s prediction:
$y_{\tau}=\begin{cases}\text{target}_{\tau}&\text{if }U(0,1)<p_{\text{tf}}(s)\text{ and training}\\ \hat{q}_{\tau}^{(5)}&\text{otherwise (median prediction)}\end{cases}$ (126)
B.5 Training Configuration
B.5.1 Weight Initialization
Model parameters are initialized using a careful scheme to ensure stable training:
$\mathbf{W}_{\text{linear}}$ $\sim\text{Xavier uniform with gain }0.01$ (127)
$\mathbf{b}$ $=\mathbf{0}$ (128)
$\mathbf{E}_{\text{embedding}}$ $\sim\mathcal{N}(0,0.01^{2})$ (129)
$\mathbf{P}_{\text{pattern}}$ $\sim\mathcal{N}(0,0.02^{2})$ (130)
B.5.2 Loss Function
The model is trained using quantile loss (pinball loss) averaged across all quantiles:
$\mathcal{L}_{\text{quantile}}$ $=\frac{1}{BH|Q|}\sum_{b=1}^{B}\sum_{\tau=1}^{H}\sum_{q\in Q}\rho_{q}(y_{b,\tau}-\hat{q}_{b,\tau}^{(q)})$ (131)
$\rho_{q}(e)$ $=\max(qe,(q-1)e)$ (132)
where $Q=\{0.05,0.1,0.25,0.4,0.5,0.6,0.75,0.9,0.95\}$ is the set of quantile levels, and $e$ is the prediction error.
B.5.3 Optimization
The model is trained using AdamW optimizer with OneCycle learning rate scheduling:
AdamW $:\beta_{1}=0.9,\beta_{2}=0.999,\text{weight decay}=0.01$ (133)
OneCycle $:\text{max\_lr}=10^{-4},\text{pct\_start}=0.05,$ (134)
$\quad\text{div\_factor}=10.0,\text{final\_div\_factor}=100.0$ (135)
Gradient clipping is applied with maximum norm 1.0, and training uses gradient accumulation every 64 steps to simulate larger effective batch sizes.
B.6 Model Scale and Efficiency
The full model configuration uses:
  • Hidden dimension: $d=1024$
  • Feed-forward dimension: $d_{\text{ffn}}=2048$
  • Number of encoder layers: $N=16$
  • Number of attention heads: $h=16$
  • Pattern memory size: $K=256$
  • Total parameters: approximately 150M
Training requires approximately 2 days on a single NVIDIA L4 GPU for one complete pass through the synthetic dataset (we only train with one epoch, since we can generate as much data as we need), while inference can be performed in real-time on standard hardware.
Appendix C Evaluation Data
After training, Mantis is evaluated on real-world forecasting tasks spanning multiple diseases, regions, and targets. For each task, the model is run out-of-the-box: no retraining, fine-tuning, or exposure to real data is performed. Forecast accuracy is assessed using both pointwise metrics such as mean absolute error (MAE), weighted interval score (WIS), and coverage metrics. Mantis’s performance is compared to established forecasting benchmarks, including the CDC Forecast Hub ensemble and individual submissions [2], as well as published models such as Chronos [18].
C.1 Primary Evaluation Tasks
C.1.1 COVID-19 Mortality Across U.S. States
To assess generalization to novel outbreak scenarios, we evaluated Mantis on weekly COVID-19 death forecasts across all 50 U.S. states, the District of Columbia, and Puerto Rico during the first two years of the pandemic (April 2020 through November 2021). This evaluation period matches the timeframe used in [2], enabling direct comparison with published CDC COVID-19 Forecast Hub results.
Data Characteristics
  • Temporal Coverage: Weekly data from April 2020 through November 2021 (87 weeks)
  • Geographic Resolution: 52 jurisdictions (50 states + DC + Puerto Rico)
  • Outcome Variable: Weekly COVID-19 deaths
  • Population Range: 580,000 (Wyoming) to 39.5 million (California)
  • Total Forecasts: Approximately 22,776 individual forecasts across all jurisdictions and horizons
  • Data Source: CDC COVID-19 surveillance data [12]
Baseline Comparisons
We benchmarked against models from the U.S. CDC COVID-19 Forecast Hub, which represented expert-curated approaches from modeling teams throughout the pandemic:
  • CDC Forecast Hub ensemble: Weighted combination of top-performing individual models
  • Individual Hub models: 28 distinct modeling approaches including statistical, mechanistic, and machine learning methods
  • Covariate utilization: Most models incorporated extensive covariate information including case counts, mobility data, and syndromic surveillance signals
  • Continuous retraining: Baselines were retrained as new data became available, reflecting operational practice
C.2 Distribution Shift Evaluation Tasks
To further assess generalization capabilities under varying degrees of distribution shift, we conducted four additional evaluations on diseases and contexts that challenge different aspects of model robustness.
C.2.1 Dengue Fever in Brazil (Vector-Borne Transmission)
We evaluated Mantis on dengue fever case forecasting across Brazilian states from 1993 to 2005, representing real-world vector-borne transmission dynamics within the model’s training distribution.
Data Characteristics
  • Temporal Coverage: Monthly data from 1993 through 2005 (156 months)
  • Geographic Resolution: Alagoas, Brazil
  • Outcome Variable: Monthly dengue fever cases
  • Population Range: Roughly 3 million
  • Total Forecasts: Approximately 4,096 individual forecasts
  • Data Source: Project Tycho [13]
  • Transmission Mode: Vector-borne (included in training simulations)
C.2.2 Scarlet Fever in the United States (Historical Context)
Scarlet fever forecasting across U.S. states from 1910 to 1950 represents historical surveillance contexts with limited infrastructure and irregular reporting patterns.
Data Characteristics
  • Temporal Coverage: Weekly data from 1910 through 1950 (40 years)
  • Geographic Resolution: 48 contiguous U.S. states (Alaska and Hawaii not yet states)
  • Outcome Variable: Weekly scarlet fever cases
  • Population Range: 200,000 to 10 million per state (1910-1950 demographics)
  • Total Forecasts: Approximately 70,560 individual forecasts
  • Data Source: Project Tycho [16]
  • Distribution Shift: Early 20th-century surveillance systems with limited medical infrastructure
C.2.3 Hepatitis B in the United States (Blood-Borne Transmission)
Hepatitis B forecasting across U.S. states from 1978 to 1988 represents chronic blood-borne infections with long latency periods—a transmission mode explicitly excluded from Mantis’s training data.
Data Characteristics
  • Temporal Coverage: Weekly data from 1978 through 1988 (11 years)
  • Geographic Resolution: 50 U.S. states
  • Outcome Variable: Weekly hepatitis B cases
  • Population Range: 400,000 to 25 million per state (1980s demographics)
  • Total Forecasts: Approximately 57,344 individual forecasts
  • Data Source: Project Tycho [14]
  • Distribution Shift: Chronic blood-borne transmission not included in training simulations
C.2.4 Smallpox in the United States (Eradicated Disease)
Smallpox forecasting across U.S. states from 1925 to 1933 represents an eradicated disease with no modern analog in the training dataset.
Data Characteristics
  • Temporal Coverage: Weekly data from 1925 through 1933 (9 years)
  • Geographic Resolution: 48 contiguous U.S. states
  • Outcome Variable: Weekly smallpox cases
  • Population Range: 500,000 to 12 million per state (1920s-1930s demographics)
  • Total Forecasts: Approximately 21,888 individual forecasts
  • Data Source: Project Tycho [15]
  • Distribution Shift: Eradicated disease with no training analog; pre-vaccine era dynamics
C.2.5 Influenza-Like Illness (ILI) in the United States (Syndromic Data)
ILI forecasting across U.S. states from 2013 to 2025 represents a challenge to Mantis due to the lack of syndromic indicators in its training set.
Data Characteristics
  • Temporal Coverage: Weekly data from 2013 through 2025 (13 years) [10]
  • Geographic Resolution: 50 U.S. states, Washington D.C., and Puerto Rico.
  • Outcome Variable: Weekly percentage of outpatient visits due to ILI.
  • Population Range: 500,000 to 40 million per state
  • Total Forecasts: Approximately 35,000 individual forecasts
  • Data Source: CMU DELPHI [11]
  • Distribution Shift: Syndromic indicator is a data type not seen during training
C.3 Baseline Model Implementations
For the distribution shift evaluation tasks, we implemented strong statistical baselines to establish performance benchmarks:
C.3.1 Naïve Persistence Model
The persistence model assumes that future values will equal the most recent observed value:
$\hat{y}_{t+h}=y_{t}\quad\text{for all forecast horizons }h$ (136)
C.3.2 Exponential Smoothing (ETS)
We implemented exponential smoothing models with optimal configuration selection via Akaike Information Criterion (AIC). For each disease and forecast window, we evaluated a comprehensive range of ETS variants:
  • Error types: Additive and multiplicative
  • Trend types: None, additive, additive damped, multiplicative, multiplicative damped
  • Seasonal types: None, additive, multiplicative
  • Training procedure: Rolling forecasts with 104 weeks (or maximum available) of historical context
  • Model selection: Best configuration selected per disease based on AIC across training period
C.3.3 Seasonal ARIMA (SARIMA)
We implemented seasonal autoregressive integrated moving average models with automatic order selection. For each disease and forecast window, the model specification was determined using the following procedure:
  • Order selection: Auto ARIMA algorithm to determine optimal $(p,d,q)\times(P,D,Q)_{s}$ parameters based on AIC
  • Seasonality: Period determined by data frequency (e.g., 52 for weekly)
  • Training procedure: Model refit at each time step using all available historical data up to the forecast origin
  • Forecast generation: Rolling forecasts with the refitted model producing multi-step-ahead predictions
  • Model selection: Best SARIMA specification selected per disease based on AIC performance during the initial training period
C.3.4 Long Short-Term Memory Network (LSTM)
We implemented a deep recurrent neural network baseline using a multi-layer LSTM architecture. The model was retrained at each forecast point to incorporate all available historical data:
  • Architecture: 6-layer LSTM with hidden dimension of 128 units per layer
  • Regularization: Dropout rate of 0.5 applied between LSTM layers
  • Data preprocessing: Standard scaler applied to normalize input time series
  • Training procedure: Model retrained from scratch at each time step using all available historical data
  • Training duration: 50 epochs per retraining with early stopping based on validation loss
  • Forecast generation: Rolling forecasts produced after retraining at each forecast origin
C.4 Data Preprocessing and Quality Control
C.4.1 Standardization
All time series were preprocessed using consistent procedures:
  • Missing value handling: Linear interpolation for isolated missing points; exclusion of series with $\geq 10\%$ missing data (this leads to most states being excluded for the historical diseases)
  • Minimum length requirements: Series with $\leq 52$ observations excluded from evaluation
C.4.2 Forecast Windows
Evaluation employed rolling forecast windows with consistent protocols:
  • Context length: Up to 112 weeks of historical data for model input
  • Forecast horizons: 4 and 8 weeks ahead for primary tasks; 8 weeks for distribution shift tasks
  • Rolling evaluation: New forecasts generated weekly/monthly as new data became available
  • Evaluation metrics: Mean Absolute Error (MAE), Mean Absolute Percentage Error (MAPE), 90% prediction interval coverage
C.5 Evaluation Scope Summary
Table 5 summarizes the complete evaluation framework, highlighting the diversity of diseases, transmission modes, geographic scales, and temporal contexts covered in our assessment.
Table 5: Summary of Mantis Evaluation Framework
Disease Location Period Frequency Transmission Distribution Shift
Influenza Michigan regions 2023-2024 Weekly Respiratory None (in-distribution)
COVID-19 U.S. states/territories 2020-2021 Weekly Respiratory Novel pathogen
Dengue Brazilian states 1993-2005 Monthly Vector-borne None (in-distribution)
Scarlet fever U.S. states 1910-1950 Annual Respiratory Historical context
Hepatitis B U.S. states 1978-1988 Annual Blood-borne Excluded transmission
Smallpox U.S. states 1925-1933 Annual Respiratory Eradicated disease
This comprehensive evaluation framework enables robust assessment of Mantis’s forecasting capabilities across the full spectrum of infectious disease scenarios, from routine seasonal diseases to novel outbreaks and historically eradicated pathogens. The combination of primary tasks (representing operational forecasting scenarios) and distribution shift challenges (testing model robustness) provides a thorough evaluation of the simulation-grounded approach’s generalization capabilities.
Appendix D Additional Results
D.1 Ablation Studies Confirm the Importance of Scale and Mechanistic Pretraining
To understand the key factors driving Mantis’s performance, we conducted two ablation studies examining the roles of architectural capacity and mechanistic pretraining in enabling accurate plug-and-play forecasting.
D.1.1 Architectural Capacity Requirements
We first evaluated whether Mantis’s forecasting capabilities could be achieved with a simpler, more compact architecture. Using the same 400 million days of synthetic training data, we replaced Mantis’s hybrid CNN-transformer architecture with the dual-branch LSTM architecture from DEFSI [9]. This LSTM model contains approximately 200 times fewer parameters than Mantis, making it significantly more lightweight and computationally efficient.
Despite identical training data and procedures, the reduced architectural capacity led to substantially degraded performance. When evaluated on COVID-19 mortality forecasting across U.S. states, the LSTM architecture achieved mean absolute error (MAE) that was 89% higher than the full Mantis model. This demonstrates that while the synthetic training paradigm is powerful, it requires sufficient architectural capacity to effectively learn and generalize from the diverse mechanistic patterns encoded in the simulation data.
D.1.2 Mechanistic Pretraining is Essential for Novel Outbreak Performance
To isolate the contribution of mechanistic pretraining, we trained the full Mantis architecture exclusively on real-world data, eliminating simulations from the training process. We evaluated on COVID-19 hospitalization forecasting. We implemented a rolling training strategy that provided the strongest possible baseline: the model was initially trained on the first few months of pandemic data, then progressively retrained with expanding datasets as more real-world observations became available. This approach mimics operational practice and ensures that the real-world-only model has access to all available historical information at each forecast point.
The results reveal the critical importance of mechanistic pretraining. When trained only on real-world COVID-19 data, forecast error increased by 262% compared to the simulation-pretrained Mantis model. This dramatic performance degradation reflects the fundamental challenge of learning outbreak dynamics from limited historical data during a novel pandemic. Real-world training data, even when accumulated over months, cannot capture the full spectrum of epidemic behaviors that might emerge. In contrast, the simulation-pretrained model draws from hundreds of millions of days of diverse outbreak scenarios, enabling it to recognize and extrapolate patterns even in unprecedented contexts.
These ablation results underscore the core insight behind Mantis: mechanistic pretraining enables models to internalize generalizable principles of epidemic dynamics that extend far beyond what can be learned from any single real-world outbreak. While architectural sophistication contributes to performance, the simulation-grounded training paradigm is the essential component that enables out-of-the-box generalization to novel diseases and outbreak contexts.
D.2 Performance Consistency Across Population Scales
To assess whether Mantis’s performance varies systematically with jurisdiction size, we analyzed the relationship between state population and forecast accuracy for the COVID-19 mortality forecasting task. We computed the mean relative MAE for each U.S. state and territory across all forecast dates in the evaluation period (April 2020 through November 2021) and examined the correlation with 2020 Census population estimates.
Figure 4 shows Mantis’s relative MAE versus state population on a logarithmic scale. We excluded Vermont from this analysis due to anomalous worse performance, perhaps due to data quality issues: in Vermont’s COVID-19 surveillance data, hospitalizations significantly lag deaths (rather than preceding them as expected), suggesting potential reporting artifacts that confound model evaluation. Among the remaining 51 jurisdictions, we observed a weak negative correlation between log population and relative MAE ($R^{2}=0.06$), meaning Mantis performs slightly better in larger states. However, the low $R^{2}$ value suggests that Mantis performance is relatively consistent across diverse population scales.
This consistency is positive given that smaller jurisdictions present distinct forecasting challenges due to noisier case trajectories and sometimes less available data. The weak relationship between population and performance suggests that Mantis’s simulation-based training enables robust forecasting across the spectrum of population sizes from states like Wyoming to California.
Figure 4: Mantis delivers consistent performance across population scales. Relative MAE versus state population for COVID-19 mortality forecasts across 51 U.S. states and territories (Vermont excluded as an outlier). Each point represents the mean relative MAE for a jurisdiction across all forecast dates from April 2020 through November 2021. Population is shown on a logarithmic scale (2020 Census). A weak negative correlation ($R^{2}=0.06$) indicates slightly better performance in larger states, but the low coefficient of determination suggests that Mantis achieves relatively uniform accuracy across diverse population sizes.